OpenStudy (anonymous):

Plsss help to solve each system of equation(a)graphically and (b) by using any algebraic method..this is a topic of system of linear equations. here`s the equations: {x-2y=8 and {x+5y=-10 Solution: (x,y)=____

OpenStudy (somethingawesome):

To solve it graphically, graph each equation and see where the lines intersect, like this: http://www.wolframalpha.com/input/?i=x-2y%3D8%2Cx%2B5y%3D-10 To solve algebraically, one method is to take one equation and solve for one of the variables, then we can plug it into the other, like this: x-2y = 8 x = 8+2y, so in the other equation, x+5y=-10 (8+2y)+5y=-10 8+7y=-10 7y=-18 y=-18/7. To find x, we can use one of those equations, or the x = 8+2y equation we found along the way: x = 8+2(-18/7) x = 20/7. So your solution ends up being (20/7, -18/7).

OpenStudy (anonymous):

so is this the final answer or is there an alternative solution?

OpenStudy (somethingawesome):

There's only one answer, but we could have used a different algebraic method. For example, we could subtract the equations: x-2y = 8 -(x+5y) = -(-10) -------------- -7y = 18 y = -18/7, and so on to find x. Both get you to the same answer, so use whichever you're more comfortable with.

OpenStudy (anonymous):

ahhh okay..thats what im thinking if i`ll use elimination by subtraction..:-D but anyway thanks for the answer and for the effort to solve this probs..God Bless u

OpenStudy (somethingawesome):

Happy to help! Good luck!